### v1

# Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up

Abstract
Large language models (LLMs) have shown remarkable performance in reasoning tasks but face limitations in mathematical and complex logical reasoning. Existing methods to improve LLMs’ logical capabilities either involve traceable or verifiable logical sequences that generate more reliable responses by constructing logical structures yet increase computational costs, or introduces rigid logic template rules, reducing flexibility. In this paper, we propose Reversal of Thought (RoT), a novel framework aimed at enhancing the logical reasoning abilities of LLMs. RoT utilizes a Preference-Guided Reverse Reasoning warm-up strategy, which integrates logical symbols for pseudocode planning through meta-cognitive mechanisms and pairwise preference self-evaluation to generate task-specific prompts solely through demonstrations, aligning with LLMs’ cognitive preferences shaped by Reinforcement Learning with Human Feedback (RLHF). Through reverse reasoning, we ultilize a Cognitive Preference Manager to assess knowledge boundaries and further expand LLMs’ reasoning capabilities by aggregating solution logic for known tasks and stylistic templates for unknown tasks. Experiments across various tasks demonstrate that RoT surpasses existing baselines in both reasoning accuracy and efficiency.
Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up
Jiahao Yuan1, Dehui Du1, Hao Zhang1, Zixiang Di1, Usman Naseem2 1East China Normal University 2Macquarie University 51275900024@stu.ecnu.edu.cn, dhdu@sei.ecnu.edu.cn, {71275902039,51265901113}@stu.ecnu.edu.cn, usman.naseem@mq.edu.au
1 Introduction
Figure 1: Comparison between CoT Yao et al. (2024); Besta et al. (2024); Yang et al. (2024a) and Reversal of Thought (RoT)
Large language models (LLMs) like Qwen Bai et al. (2023), Llama Dubey et al. (2024), and GPT-4 Achiam et al. (2023) have demonstrated remarkable performance in various reasoning tasks via single-step prompting with few shots upon scaling model size Plaat et al. (2024) but remain restricted in mathematical and intricate logical reasoning domains Arkoudas (2023); Stechly et al. , which has spurred more effective multi-step Chain-of-Thought (CoT) prompting Wei et al. (2022) approaches for activating step-by-step logical capabilities. However, LLMs are prone to unfaithfulness, resulting in cascaded intermediate errors Bao et al. (2024); Yang et al. (2024b).
Recent studies have advanced CoT to guide LLMs, mainly through either multi-step prompting such as introducing planning-and-solve Plaat et al. (2024); Yang et al. (2024a), self-consistency Narang et al. and recursive reasoning process Lee and Kim (2023); Yu et al. (2024) through Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024), or multi-role Zhang et al. ; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased the computational cost of inference or API access cost due to the multi-step inference. Buffer-of-Thought (BoT) Yang et al. (2024a) attempts to reduce thinking steps by leveraging Retrieval-Augmented Generation (RAG) to retrieve gold thought templates from the buffer. However, it sacrifices flexibility due to the initialization of pre-set manual thought templates. Therefore, achieving accurate reasoning in LLMs while minimizing resource consumption remains a significant challenge.
In summary, existing methods primarily rely on multi-query CoT which injects knowledge Suzgun and Kalai (2024); Plaat et al. (2024) or data structure Yao et al. (2024); Besta et al. (2024) to optimize decisions making, and encounter three significant limitations: (1) limitation in logical reasoning: Despite attempting different logic data structures Yao et al. (2024); Besta et al. (2024); Yang et al. (2024a), an effective initiative Chain-of-Thought paradigm that suits and improves logical reasoning remains elusive Bao et al. (2024); (2) unfaithfulness and cascaded errors: Single-step or multi-step methods are liable to cause LLMs to output hallucinations, leading to cascading logic errors Bao et al. (2024); (3) Trade-off between enhanced logic capabilities and resource consumption: Recent CoT advancements via multi-step or multi-role prompting increase costs and achieving a balance between logical flexibility, accuracy, and cost is of great significance for practical application.
To address above limitations, we propose Reversal of Thought (RoT), a novel framework that enables LLMs to explore cognitive preference on logical pseudocode solely using reverse prompting with given demos without additional task-related affirmations, as depicted in Figure 1. Our key contributions are as follows:
  • To the best of our knowledge, we are the first to introduce a reversal reasoning for cognitive preference that enhances logical reasoning in LLMs by combining meta-cognitive with cognitive preference, resulting in a more cost-efficient and error-resilient framework for complex tasks.
  • We propose a Preference-Guided Reverse Reasoning framework that enhances LLMs’ task cognition by employing a reverse reasoning warm-up strategy and preference-based self-evaluation to improve logical reasoning based on LLMs’ cognitive preferences.
  • We introduce logical symbols in pseudocode for algorithmic planning and problem-solving, guided by meta-cognitive mechanisms, which improves the LLMs’ capacity for structured and accurate logical reasoning.
  • We introduce a Cognitive Preference Manager to evaluate knowledge boundaries, enabling the automatic adaptation of cognitive preference styles for unknown logic tasks and efficient aggregation of solution logic for known tasks.
2 Related Work
2.1 Chain-of-Thought (CoT) Prompting
Chain-of-Thought (CoT) prompting Wei et al. (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic. Recent studies primarily aimed at improving logical accuracy by introducing more external validation such as self-consistency Narang et al. ; Yu et al. (2024) or more hierarchical information such as Least-to-Most Zhou et al. , Cumulative-Reasoning Zhang et al. and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors or poor flexibility. Additionally, numerous studies also proposed more standardized recursive or backtracking branch forms from the logical data structure, including Tree-of-Thought (ToT) Yao et al. (2024), Graph-of-Thought (GoT) Besta et al. (2024) and Buffer-of-Thought (BoT) Yang et al. (2024a). However, an efficient logical reasoning method that strikes a balance among logical accuracy, flexibility, and cost has yet to be discovered. Our method is activated through meta cognition Fleur et al. (2021) by introducing reverse reasoning to form effective LLMs-taste prompts within cognitive preference Uddin (2021) for plan-and-solve with logical pseudocode at least.
2.2 Knowledge Boundary for Enhancing Large Language Models
Integrating knowledge boundary within LLMs has emerged as a prospective strategy for enhancing their ability to avoid reasoning hallucinations of unknown knowledge through knowledge boundary constraints which requires additional algorithmic efforts Yin et al. (2024), external graph knowledge Tian et al. (2024), and training consumption Sun et al. (2024). Additionally, they focus on avoiding responses to unknown or incorrect prompts rather than proposing bold and proactive solutions to expand knowledge boundary in a heuristics without training.
Figure 2: Architecture of Reversal-of-Thought (RoT). RoT comprises two primary components: Preference Guided Reverse Reasoning, which enhances logical reasoning by activating LLMs’ cognitive preferences, and Cognitive Preference Manager, which assesses knowledge boundaries and adapts cognitive styles for various tasks.
We proposed a prompt-based method utilizing LLMs pretrained knowledge boundary, inspired by meta cognition Fleur et al. (2021) and cognitive preference for unknown knowledge Uddin (2021). Our method conducts reverse prompting on probing knowledge through demonstrations to obtain LLMs-taste problem cognitions, aggregates and distills original prompt into cognitive preference version.
3 Reversal of Thought
3.1 Overview
Tell me and I forget. Teach me and I remember. Involve me and I learn.
Franklin (2005)
As the aforementioned wisdom related to human cognitive learning implies, merely telling or teaching is inadequate Bao et al. (2024). Moreover, most LLMs have undergone extensive pre-training Achiam et al. (2023); Bai et al. (2023); Dubey et al. (2024) and reinforcement learning from human feedback (RLHF) Ouyang et al. (2022), instilling in LLMs a propensity for specific cognitive patterns, which manifests in two progressive layers of LLMs-taste description: (1) Stylistic template: encompassing grammatical and syntactic structures in descriptions. (2) Solution logic: comprising problem-solving reasoning and methodological cues. Therefore, Reversal of Thought (RoT) involves answering the following two research questions (RQs):
  • RQ1: How to make LLMs output preference cognitive templates and logic for specific tasks and activate known cognitive boundaries?
  • RQ2: How to autonomously use cognitive templates with incorrect response to expand the possible knowledge boundaries?
To activate and enhance LLMs logical flexibility, accuracy, and the ability to autonomously construct meta-cognition without training for logical reasoning, inspired by meta-cognition Fleur et al. (2021) and cognitive preference Uddin (2021); Zhou et al. (2023); Margatina et al. (2023), we introduce Reversal of Thought (RoT), a cost-effective paradigm that enables LLMs to first explore cognitive preference on logical pseudocode solely through given examples without additional task-oriented affirmation, activates the pre-trained known logic under Reverse Reasoning Warm-up (detailed in section 3.2), and then optimizes the original prompt for LLMs-taste prompt via Cognitive Preference Manager (detailed in section 3.2) to determine the transfer of cognitive preference style for unknown logic template and aggregation of known solution logic, as depicted in Figure 2.
Algorithm 1 Preference-Guided Reverse Reasoning (PGRR)
0:  $P$: Initial prompt, $D$: Input-output demonstrations, $warm$: Number of warm iterations
1:  (1) Reverse Reasoning Warm-up:
2:  for $i=1$ to $warm$ do
3:     $R^{(i)}\leftarrow\mathcal{M}_{\text{LLM}}(P_{r},D,i)$Generate candidate responses}
4:     $P_{res}^{i}\leftarrow\frac{1}{|R^{(i)}|}\sum_{R_{i,j}\in R^{(i)}}\exp(P(R_{i,j}|P_{r},D))$
5:  end for
6:  $R\leftarrow\bigcup_{i=1}^{warm}R^{(i)}$Collect all responses} (2) Pairwise Preference Evaluation:
7:  for $i=0$ to $warm-1$ do
8:     $P_{pre}(R_{i+1}\succ R_{i})\leftarrow\exp(\mathcal{M}_{\text{LLM}}(P_{eval},R_{i+1},R_{i}))$
9:  end for
10:  for $i=0$ to $warm-1$ do
11:     for $j=0$ to $i-1$ do
12:        $P_{pre}(R_{i}\succ R_{j})\leftarrow P_{pre}(R_{i}\succ R_{j})$Utilize preference transitivity}
13:     end for
14:  end for(3) Preference-Guided Ranking:
15:  for $i=1$ to $warm$ do
16:     $\bar{P}_{pre}(R_{i})\leftarrow\frac{1}{warm-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{warm}P_{pre}(R_{i}\succ R_{j})$
17:  end for
18:  $P_{opt}\leftarrow\arg\max_{R_{i}}\left(\frac{P_{res}^{i}+\bar{P}_{pre}(R_{i})}{2}\right)$
19:  return  $P_{opt}$Optimal LLMs-taste prompt}
3.2 Reverse Reasoning with Meta-cognition
Preference-Guided Reverse Reasoning.
Inspired by RLHF Ouyang et al. (2022) utilizing preference data, and to derive high-cognitive preference prompt $P^{*}$ that enhance logical reasoning in LLMs, we propose a Preference-Guided Reverse Reasoning (PGRR) framework (detailed in alogrithm 1) mapping input-output demonstrations $D$ from an initial prompt $P$ to an optimal LLM-taste prompt $P_{opt}$.
(1) Reverse Reasoning Warm-up.
We query the LLM $\mathcal{M}_{\text{LLM}}$ with a reversal prompt and demonstrations $\{P_{r},D\}$ (detailed in figure 3) for $warm$ iterations, generating a set of candidate cognitive preference responses $R=\{R_{1},R_{2},\dots,R_{warm}\}$ and their corresponding average probabilities $P_{res}^{i}$:
$R=\bigcup_{i=1}^{\text{warm}}R^{(i)}=\bigcup_{i=1}^{\text{warm}}\mathcal{M_{LLM}}(P_{r},D,i),$ (1)
$P_{res}^{i}=\frac{1}{|R_{i}|}\sum_{R_{i,j}\in R}\exp(P(R_{ij}|P_{r},D)).$ (2)
where $R^{(i)}$ represents the $i$-th generated response. $\mathcal{M}_{\text{LLM}}(P_{r},D,i)$ is the model output based on the reversal prompt $P_{r}$ and demonstrations $D$ for the $i$-th iteration. $P(R_{ij}|P_{r},D)$ denotes log probability for each token $R_{ij}\in R_{i}$ from LLMs.
(2) Pairwise Preference Evaluation.
To acquire the most LLMs-taste prompt, we pair candidate responses $R$ as data pairs $\{R_{i},R_{i+1}\}$ where $i=0,1,\cdots,warm-1$ to calculate the relative preference $P(R_{i+1}\succ R_{i})$ through LLM’s self-evaluation of its preference for $R_{i+1}$ over $R_{i}$, formally define as:
$P_{pre}(R_{i+1}\succ R_{i})=\exp(\mathcal{M_{LLM}}(P_{eval},R_{i+1},R_{i}))$ (3)
where $\mathcal{M_{LLM}}(Prompt_{eval},R_{i+1},R_{i})$ represents that require LLM to select more preferred data through $P_{eval}$ with a structure as "Please choose your more preferred instruction (A/B): (A).$R_{i+1}$; (B).$R_{i}$".
Following the principle of preference transitivity Liu et al. (2024), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
$P_{pre}(R_{i}\succ R_{j})=\begin{cases}1&i=j\\ \prod_{k=j}^{i-1}P_{pre}(R_{k+1}\succ R_{k})&i>j\\ 1-{P_{pre}(R_{j}\succ R_{i})}&i<j\end{cases}$ (4)
(3) Preference-Guided Ranking.
To identify the most LLMs-preferred and high-quality response, we compute each response $R_{i}$’s overall preference score $\bar{P}_{pre}(R_{i})$, and averaging both average probabilities $P_{res}^{i}$ in matrix $P_{pre}$ and preference score $\bar{P}_{pre}(R_{i})$ to obtain the best LLM-taste prompt $P_{opt}$:
$\bar{P}_{pre}(R_{i})$ $=\frac{1}{warm-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{warm}P_{pre}(R_{i}\succ R_{j}),$ (5)
$P_{opt}$ $=\arg\max_{R_{i}}\left(\frac{P_{res}^{i}+\bar{P}_{pre}(R_{i})}{2}\right).$ (6)
Reverse Logic for Meta-cognition.
Within reverse reasoning, we further follow meta-cognitive Suzgun and Kalai (2024) using plan-and-solve by integrating logical algorithm pseudo-code to improve reasoning comprehension. And we incorporate fundamental mathematical logic symbols, including logical operators, quantifiers, inequalities and conditional statements, to facilitate model reasoning detailed in Figure 3.
Prompt for Reverse Reasoning Instruction
You are a highly distinguished expert in mathematics and information reasoning. Based on the given example, define the specific task, including the task definition, pseudocode, logical pseudocode, case examples, and input-output format.
1. Understand Task Description:
Meticulously study demonstrations to deeply understand generic task description.
2. Plan Generic Pseudocode:
Provide pseudocode in text form and plan an efficient algorithm to complete the task with your experiences.
3. Formulate Logical Pseudocode:
Convert the pseudocode into generic logical algorithm pseudocode using ONLY logical symbols:
Logical Operators:
Conjunction: $A\land B$ ; Disjunction: $A\lor B$
equivalence: $A\equiv B$ , Negation: $\neg A$
Quantifiers:
Universal quantifier: $\forall x$ ; Existential quantifier: $\exists x$
Inequalities:
Less than: $x<y$ ; Greater than: $x>y$
Less than or equal to: $x\leq y$
Greater than or equal to: $x\geq y$
Equals: $x=y$ ; Not equals: $x\neq y$
Conditional Statements:
If $A$ then $B$: $A\supset B$
If $A\land B$ then $C$: $(A\land B)\supset C$
If $A\lor B$ then $C$: $(A\lor B)\supset C$
If $\forall x(P(x))$ then $Q$: $\forall x(P(x))\supset Q$
If $\exists x(P(x))$ then $Q$: $\exists x(P(x))\supset Q$ etc.
Input: [Demonstration] Output: [Output]
Figure 3: Prompt for Reverse Reasoning
3.3 Cognitive Preference Manager
Algorithm 2 Cognitive Preference Manager (CPM)
0:  $P$: Original prompt, $P^{*}$: Reverse-reasoned prompt, $\mathcal{M}_{emb}$: Offline LLM embedding model, $\delta$: Similarity threshold
1:  $P_{task}\leftarrow P,P^{*}_{task}\leftarrow P^{*}$
2:  $s\leftarrow\mathcal{M}_{emb}(P_{task},P^{*}_{task})$ Calculate embedding similarity between $P_{\text{task}}$ and $P^{*}_{\text{task}}$.
3:  if $s\geq\delta$ then
4:     Known detected: Enhance and refine $P$ to optimized instructions $P_{final}$.
5:     Aggregate relevant task-specific knowledge.
6:     return  $P_{final}$.
7:  else
8:     Unknown detected: Adapt and expand $P_{task}$.
9:     Leverage cognitive preference templates $T$ and $P$ to generate optimized instructions $P_{final}$.
10:     return  $P_{final}$.
11:  end if
Cognitive Preference Manager.
After reverse reasoning for LLMs-cognitive description $P^{*}$, We introduce an offline-deployed LLM embedding model $\mathcal{M}_{emb}$ to assist Cognitive Preference Manager (CPM) in determining whether reverse reasoning under reverse prompt $P_{r}$ and demonstrations $D$ reaches the knowledge boundary or cognitive error by calculating the similarity and setting a threshold $\delta$ (0.6$\sim$0.8 is recommended for optimal performance in distinguishing knowledge boundaries) between orginal task defination $P_{task}$ from $P$ and LLMs-cognitive task defination $P^{*}_{task}$ from $P^{*}$, and finally get a cognitive signal $S_{cog}$, formally:
$S_{\text{cog}}=\begin{cases}unknown&,\texttt{sim}\left(\mathcal{M}_{\text{emb}}(P_{\text{task}}),\mathcal{M}_{\text{emb}}(P^{*}_{\text{task}})\right)<\delta\\ known&,\texttt{sim}\left(\mathcal{M}_{\text{emb}}(P_{\text{task}}),\mathcal{M}_{\text{emb}}(P^{*}_{\text{task}})\right)\geq\delta\end{cases}$ (7)
where $\texttt{sim}(*)$ is a cosine similarity function that computes the similarity between two embedding vectors.
By efficiently evaluating cognitive results, CPM integrates alternative aggregation strategies for $\mathcal{M_{LLM}}$ based on $S_{cog}$ 111we use $S_{cog}$ to conduct reverse evaluation of GPT-4 in our experiments., as detailed in appendix A: (1) Solution logic aggregation for known tasks: $\mathcal{M_{LLM}}$ merges beneficial aspects from the original prompt $P$ with the LLM-taste prompt $P^{*}$ to create the final prompt $P_{\textit{final}}$. (2) Stylistic template aggregation for unknown $\mathcal{M_{LLM}}$ extracts a cognitive preference template $T$ from the incorrect context in the LLM-taste prompt, and integrates meta-cognitive elements from the original prompt $P$ into $T$ to construct the final prompt $P_{final}$.
Consequently, we utilize the final LLM-preferred prompt to query the LLM $\mathcal{M_{LLM}}$ with a specific problem input, $problem$, to obtain the final logical answer, $answer$:
$answer=\mathcal{M_{LLM}}(P_{final},problem)$ (8)
4 Experiments
4.1 Datasets and Tasks
To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al. (2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al. (2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al. (2024) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24. (2) BIG-Bench Suzgun et al. (2023); Srivastava et al. (2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting from BIG-Bench Hard Suzgun et al. (2023) and Checkmate-in-One from BIG-Bench suite Srivastava et al. (2023); (3) Python Puzzles Schuster et al. (2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels.;(4) Multilingual Grade School Math (MGSM) Shi et al. (2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al. (2021), featuring translations of a subset of examples into ten diverse languages.(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three specified words verbatim.
4.2 Baselines
In our experiments, we compare RoT with five classic and latest state-of-the-art prompting baselines:
  • CoT Prompting: Following Suzgun and Kalai (2024); Yang et al. (2024a), we employ GPT-4 to decompose instruction into logic intermediate reasoning steps activated by "Let’s think step by step".
  • Meta-Prompting: Suzgun and Kalai (2024) introduced general, task-agnostic prompts as a scaffold to guide LLMs effectively perform logic tasks.
  • Graph-of-Thought (GoT): Besta et al. (2024) modeled reasoning as a graph, where nodes are thoughts and edges define their relationships to solve complex problems.
  • Tree-of-Thought (ToT): Yao et al. (2024) organized reasoning in a tree structure, allowing LLMs to explore multiple thought paths and select the most promising ones for problem-solving, enhancing their complexity management in logical reasoning tasks.
  • Buffer of Thought (BoT): Yang et al. (2024a) introduced a meta-buffer that stores high-level thought templates, allowing for the adaptive retrieval and instantiation of relevant templates to enhance reasoning efficiency.
4.3 Experiment Setup
To ensure fair comparisons with previous methods, we utilize GPT-4, the latest model supporting $logprobs$  [^2] https://platform.openai.com/docs/api-reference/introduction, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study. For the $warm$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API. Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency. For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^3] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al. (2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
4.4 Evaluation Metrics
Task CoT (NeurIPS) Wei et al. (2022) Meta-prompting Suzgun and Kalai (2024) ToT (NeurIPS) Yao et al. (2024) GoT (AAAI) Besta et al. (2024) BoT (NeurIPS) Yang et al. (2024a) RoT (Ours)
1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot
Game of 24 31.5 33.2 65.4 66.8 73.3 74.1 73.0 73.5 82.8 82.5 97.5 97.8
Geometric Shapes 57.3 60.8 76.8 78.3 58.8 60.1 56.1 57.8 90.1 90.9 94.8 95.0
Multi-Step Arithmetic 83.4 85.3 89.5 90.6 88.5 90.3 87.5 88.8 94.6 96.8 98.6 99.0
WordSorting 80.8 82.5 97.8 98.6 95.0 95.8 95.6 96.8 99.2 99.6 99.8 100.0
Checkmate-in-One 35.8 36.5 58.0 57.5 49.8 48.8 49.0 51.2 87.2 88.1 91.8 92.0
Python Puzzles 35.3 35.8 45.0 45.5 43.8 45.1 42.5 43.5 51.8 52.6 54.8 57.5
MGSM (avg) 83.8 84.2 85.2 85.4 84.5 86.1 85.8 86.8 87.5 87.3 88.9 90.1
Sonnet Writing 66.5 70.2 78.8 79.8 68.9 69.8 68.0 62.5 79.0 79.5 89.5 92.0
Table 1: Comparison of RoT with baselines across 1-shot and 2-shot settings for reasoning accuracy.
Knowledge Boundary and Cognitive Preference Consistency.
As described in Section 3.2, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$. We first evaluate GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^4] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot. and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
Reasoning Accuracy.
Following Suzgun and Kalai (2024); Yao et al. (2024), we introduce a LLM555we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong). We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
Reasoning Efficiency.
We evaluate reasoning efficiency by calculating ${t_{eff}}$, calculated as the average time spent per task across all samples in the test dataset:
$t_{eff}=\frac{1}{N}\sum_{j=1}^{M}\sum_{i=1}^{N}t_{ij}$ (9)
where $N$ is the total number of tasks, $M$ is the total number of samples, and $t_{ij}$ represents the time taken for the $i$-th task on the $j$-th sample.
5 Results and Discussion
Task 1-shot 2-shot
Kno. Unkno. Kno. Unkno.
Game of 24
MGSM (avg)
Multi-Step Arithmetic
WordSorting
Python Puzzles
Geometric Shapes
Checkmate-in-One
Sonnet Writing
Table 2: Knowledge Boundary under Reversal Reasoning. Kno. for Known, Unkno. for Unknown.
5.1 Knowledge Boundary and Cognitive Preference Consistency
Knowledge Boundary.
As shown in Table 2, our findings provide preliminary support for our hypothesis concerning cognitive knowledge boundaries of LLMs in reverse reasoning across current task benchmarks in one-shot and two-shot settings Yao et al. (2024); Suzgun et al. (2023); Srivastava et al. (2023); Schuster et al. (2021); Shi et al. (2022); Suzgun and Kalai (2024). Notably, GPT-4 excels in structured reasoning tasks such as Game of 24, Geometric Shapes, and Checkmate-in-One, consistently achieving strong results in both one-shot and two-shot settings. In contrast, tasks such as MGSM (avg) and Python Puzzles fall into an unknown category in the one-shot scenario, stemming from multi-source problems that are challenging to make reverse reasoning without sufficient context, indirectly supporting the rationale for our CPM approach, detailed in appendix B.
Cognitive Preference Consistency.
Moreover, our human evaluation of $\textit{Con}_{cog}$ produced a high score of 4.32 in the two-shot setting and 4.01 in the one-shot setting, further validating that LLMs exhibit cognitive preferences shaped by RLHF Ouyang et al. (2022).
5.2 Reasoning Accuracy and Efficiency
Reasoning Accuracy.
As shown table 1, our RoT model consistently outperforms all baseline models across various logic reasoning tasks. Notably, compared with the best BoT, we observe significant accuracy improvements in 1-shot and 2-shot settings, particularly evidenced by increases of 14.7% and 15.3% on the Game of 24, 4.7% and 4.1% on Geometric Shapes, and 4.6% and 4.5% on Checkmate-in-One, demonstrating our flexibility and efficiency of leveraging cognitive preference in LLMs to activate logic capabilities through reverse reasoning, particularly with cognitive preference management.
Reasoning Efficiency.
Figure 4: Inference time comparison
Dataset RoT w/o PGRR w/o Logic w/o CPM
1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot
Game of 24 97.5 97.8 77.5 77.8 86.0 86.5 92.0 93.5
Geometric Shapes 94.8 95.0 74.8 75.0 83.0 83.5 90.0 90.5
Multi-Step Arithmetic 98.6 99.0 78.6 79.0 87.5 88.0 95.0 95.5
WordSorting 99.8 100.0 79.1 81.0 89.0 89.2 98.0 98.5
Checkmate-in-One 91.8 92.0 80.8 82.0 81.0 83.5 98.0 88.5
$\times$×Python Puzzles 54.8 57.5 44.3 47.1 50.0 50.5 41.8 52.5
$\times$×MGSM (avg) 88.9 90.1 85.8 86.1 86.0 86.5 85.2 86.1
Sonnet Writing 89.5 92.0 72.5 73.0 87.0 77.5 89.0 91.5
Table 3: Ablation study of RoT across various tasks in 1-shot and 2-shot settings. indicates known tasks in both 1-shot and 2-shot; $\times$ indicates unknown tasks in 1-shot.
As shown in Figure 4, RoT achieves competitive performance in reasoning efficiency, outperforming baselines and being second only to BoT, while BoT requires numerous initial golden thought templates in advance, which subtly limits its flexibility. In contrast, RoT focuses on exploring LLM-taste prompts, requiring only data for reasoning in subsequent steps after reverse reasoning warm-up.
5.3 Ablation Study
As shown in Table 3, we conducted three ablation studies to evaluate key components: (1) w/o PGRR: Removing Preference-Guided Reverse Reasoning(PGRR) for exploring LLMs-taste prompts; (2) w/o Logic: Excluding mathematical logic for pseudo-code plan-and-solve; and (3) w/o CPM: Eliminating Cognitive Preference Manage(CPM) for both known and unknown tasks.
Impact of PGRR.
Excluding Preference-Guided Reverse Reasoning (PGRR) for exploring LLMs-taste prompts regarding tasks results in a significant reduction in overall task performance, with decreases of up to 20% in tasks like the Game of 24 (97.8% to 77.8%) and WordSorting (100% to 81%), indicating w/o PGRR weaken model’s cognition to task-specific requirements.
Impact of Logic.
Removing mathematical logic from RoT leads to notable declines in tasks requiring structured problem-solving, such as Multi-Step Arithmetic (99.0% to 88.0%) and Checkmate-in-One (92.0% to 83.5%), underscoring w/o Logic negatively affects structured problem-solving for complex reasoning.
Impact of CPM.
Lacking Cognitive Preference Manager (CPM) has a particularly pronounced effect on unknown tasks with Python Puzzles dropping from 57.5% to 52.5% and MGSM from 90.1% to 86.1%, indicating w/o CPM weaken RoT’s flexibility for tackling known tasks (case study for knwon task detailed in appendix B.1) and unknwon tasks (case study for unknwon task is detailed in Appendix B.2).
6 Conclusion
In this paper, we propose Reversal of Thought (RoT), a novel framework to enhance the logical reasoning capabilities of LLMs. By integrating reverse reasoning with meta-cognitive mechanisms and cognitive preference management, RoT improves reasoning accuracy and efficiency while minimizing computational costs, which leverages Preference-Guided Reverse Reasoning and Cognitive Preference Manager, which optimally aligns LLM reasoning processes with their cognitive preferences shaped by their pretraining and RLHF. Comprehensive experiments across diverse reasoning tasks demonstrate that RoT consistently outperforms state-of-the-art baselines in both known and unknown task scenarios, demonstrating the potential to expand knowledge boundaries through cognitive preference template. Our research provides valuable insights into future studies focused on further enhancing LLMs’ reasoning capacities by dynamically exploring cognitive preferences for complex reasoning tasks.
Limitations
Reversal of Thought (RoT) introduce a reverse reasoning warm-up to activate cognitive preference for LLMs to enhance logic capabilities and introduce a cognitive preference manager to determine knowledge boundary and utilize cognitive preference for known and unknown tasks.
While RoT has performed exceptionally in logic accuracy and efficiency, We discuss major challenge in its reliance on two-shot demonstration inputs involving two distinct problem cases. We observed that RoT may struggles with one-shot learning in multi-source tasks. we partially and effectively mitigates this issue through the integration of Cognitive Preference Manager (CPM) and two-shot learning.
In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks.
Ethical Considerations
Since the datasets used in our experiments focus on pure mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity, which are all sourced from publicly available datasets Yao et al. (2024); Suzgun et al. (2023); Srivastava et al. (2023); Schuster et al. (2021); Shi et al. (2022); Suzgun and Kalai (2024) and devoid of any personal privacy or sensitive ethical information. Therefore, we do not identify any immediate ethical concerns regarding our current work. Additionally, we conduct human evaluations of cognitive preference consistency with three anonymous professional annotators following our instructions (detailed in appendix C), and they receive equivalent compensation.
Appendix A Prompt for CPM (Known/Unknown)
Prompt for CPM (Known/Unknown) Instruction
You are an expert in information synthesis, proficient in combining complementary insights and extracting essential details from the viewpoints of the distilled task definition, detailed generic logical pseudocode, case example, and input-output format.
1. For Known:
The Reversal Prompt and Benchmark Prompt should complement each other. Analyze the content and structure of both prompts to identify their accuracy, similarities and differences. Synthesize the key points and integrate them into a unified and coherent output.
2. For Unknown:
Extract a cognitive preference template $T$ from any inaccuracies in the LLM-taste prompt. Integrate meta-cognitive elements from the original prompt $P$ into this template to enhance $T$.
Input:
LLM-Taste Prompt:[LLM-Taste Prompt]
Benchmark Prompt:[Benchmark Prompt]
Output: [Output]
Appendix B Case Study
B.1 Case Study for Known Task (Game of 24)
$P_{final}$ Case study for known task after CPM Task Defination
The task is to find a feasible mathematical expression using the four input numbers (a, b, c, d) in the order they are given, such that the result equals 24. The feasible solution should involve basic arithmetic operations (+, -, $\times$, /) and parentheses if necessary.
Logical Pseudocode
$\forall P(A,B,C,D)$ in permutations of [A, B, C, D]. $\forall op_{n}$ in (+, -, $\times$, $\div$) where $n\in\{0,1,2,3\}$. For each permute $\neg\exists(Expression==24)\lor(Expression=permute[0]\;op_{1}\;permute[1]\\ ;op_{2}\;permute[2]\;op_{3}\;permute[3])$: $\exists(Expression==24)\supset Print(Expression)\land Terminate\;Process.$ Print("No feasible solution exists."). Case Examples: Example 1: Input: 4 6 7 1 Output: One feasible solution (e.g., $6/(1-(4/7))=24$). Example 2: Input: 4 2 2 1 Output: "No feasible solution exists." Input-Output Format:
Input: Four integers separated by space (e.g., "3 3 8 8").
Output: If a feasible solution exists, output a string indicating one possible solution (e.g., "One feasible solution (e.g., $8\div(3-8\div 3)=24$)."). If no solution exists, output the string "No feasible solution exists."
B.2 Case Study for Unknown Task (MGSM)
$P_{final}$ Case study for unknown task after CPM Task Defination
The task is to solve a math problem text with a description of a situation. Your task is to calculate the answer and provide it in the format specified.
Logical Pseudocode
Input $u$ from the user in natural language. Parse $u$ for numbers and contextual information. Extract numbers $\alpha$, $\beta$, $\gamma$, etc. Identify contextual clues and operations, such as addition (+), subtraction (-), multiplication (*), division (/), and other implicit operations (e.g., percentages, halves, totals). For each identified operation: If operation is related to percentages, interpret $\times$ or division as needed. If operation is additive (+), perform $Add(\alpha,\beta)$. If operation is subtractive (-), perform $Subtract(\alpha,\beta)$. If operation is multiplicative (*), perform $Multiply(\alpha,\beta)$. If operation is divisive (/), perform $Divide(\alpha,\beta)$. If implicit operations (e.g., "half that much"), interpret accordingly. Handle complex structures, such as total amounts or remainders: Use context to evaluate remaining quantities (e.g., after consumption, sales). Apply operations in sequential order based on context. Calculate the final result based on all interpreted operations and numbers. Return the final result as a natural number or in currency (if applicable). Case Examples: Example 1:
Input: A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?
Output: 3
Input-Output Format:
Input: A natural language string describing a scenario that involves certain arithmetic operations.
Output: A single integer or float number that is the result of the operations described in the scenario.
Appendix C Instructions for Human Evaluation.
To evaluate the cognitive preference consistency $Con_{cog}$ between LLMs-taste prompt for task cognitions across different tasks, annotators adhere to the following guidelines:
Criteria.
(1) Style Consistency: How well responses maintain a consistent stylistic tone across different tasks. (2) Coherence: Internal logic and smooth flow of the response.
Scoring.
Responses are evaluated on a scale from 0 to 5. A score of 4-5 signifies exemplary style consistency, while 3-4 indicates strong consistency with minor variations. Scores of 2-3 represent adequate style but with noticeable inconsistencies, and 1-2 denote inconsistent style across tasks. A score of 0-1 reflects complete inconsistency.

### v2

@@
 Existing methods to improve LLMs’ logical capabilities either involve traceable or verifiable logical sequences that generate more reliable responses by constructing logical structures yet increase computational costs, or introduces rigid logic template rules, reducing flexibility.
-In this paper, we propose Reversal of Thought (RoT), a novel framework aimed at enhancing the logical reasoning abilities of LLMs.
-RoT utilizes a Preference-Guided Reverse Reasoning warm-up strategy, which integrates logical symbols for pseudocode planning through meta-cognitive mechanisms and pairwise preference self-evaluation to generate task-specific prompts solely through demonstrations, aligning with LLMs’ cognitive preferences shaped by Reinforcement Learning with Human Feedback (RLHF).
-Through reverse reasoning, we ultilize a Cognitive Preference Manager to assess knowledge boundaries and further expand LLMs’ reasoning capabilities by aggregating solution logic for known tasks and stylistic templates for unknown tasks.
+In this paper, we propose Reversal of Thought (RoT), a plug-and-play and cost-effective reasoning framework designed to enhance the logical reasoning abilities of LLMs during the warm-up phase prior to batch inference.
+RoT utilizes a Preference-Guided Reverse Reasoning warm-up strategy, which integrates logical symbols for pseudocode planning through meta-cognitive mechanisms and pairwise preference self-evaluation to generate task-specific prompts solely through demonstrations, aligning with LLMs’ cognitive preferences shaped by RLHF.
+Through reverse reasoning, we utilize a Cognitive Preference Manager to assess knowledge boundaries and further expand LLMs’ reasoning capabilities by aggregating solution logic for known tasks and stylistic templates for unknown tasks.
 Experiments across various tasks demonstrate that RoT surpasses existing baselines in both reasoning accuracy and efficiency.
+[…]
+[ roundcorner=5pt, backgroundcolor=UserExampleBg, linecolor=UserExampleTitle, outerlinewidth=0.5pt, frametitlebackgroundcolor=UserExampleTitle, frametitlefont=, ]user_example
 Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up
@@
 (2024a), self-consistency Narang et al.
-and recursive reasoning process Lee and Kim (2023); Yu et al.
+; Wang et al.
+(2024) and recursive reasoning process Lee and Kim (2023); Yu et al.
 (2024) through Tree-of-Thought (ToT) Yao et al.
@@
 (2024), or multi-role Zhang et al.
-; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased the computational cost of inference or API access cost due to the multi-step inference.
+; Suzgun and Kalai (2024) to enhance logical capabilities and mitigate hallucination, yet this has stealthily increased inference cost due to the multi-step inference.
 Buffer-of-Thought (BoT) Yang et al.
@@
 (2024); (3) Trade-off between enhanced logic capabilities and resource consumption: Recent CoT advancements via multi-step or multi-role prompting increase costs and achieving a balance between logical flexibility, accuracy, and cost is of great significance for practical application.
-To address above limitations, we propose Reversal of Thought (RoT), a novel framework that enables LLMs to explore cognitive preference on logical pseudocode solely using reverse prompting with given demos without additional task-related affirmations, as depicted in Figure 1.
+To address above limitations, inspired by meta-cognition Fleur et al.
+(2021) and and cognitive preference Uddin (2021); Zhou et al.
+(2023); Margatina et al.
+(2023), we propose Reversal of Thought (RoT), a plug-and-play and cost-effective framework that enables LLMs to explore cognitive preference on logical pseudocode solely using reverse prompting with given demos without additional task-related affirmations, as depicted in Figure 1.
 Our key contributions are as follows:
-  • To the best of our knowledge, we are the first to introduce a reversal reasoning for cognitive preference that enhances logical reasoning in LLMs by combining meta-cognitive with cognitive preference, resulting in a more cost-efficient and error-resilient framework for complex tasks.
+  • To the best of our knowledge, we are the first to introduce a reversal reasoning for cognitive preference that enhances logical reasoning in LLMs by combining meta-cognitive with cognitive preference, resulting in a more modular and cost-efficient framework for complex tasks.
   • We propose a Preference-Guided Reverse Reasoning framework that enhances LLMs’ task cognition by employing a reverse reasoning warm-up strategy and preference-based self-evaluation to improve logical reasoning based on LLMs’ cognitive preferences.
-  • We introduce logical symbols in pseudocode for algorithmic planning and problem-solving, guided by meta-cognitive mechanisms, which improves the LLMs’ capacity for structured and accurate logical reasoning.
   • We introduce a Cognitive Preference Manager to evaluate knowledge boundaries, enabling the automatic adaptation of cognitive preference styles for unknown logic tasks and efficient aggregation of solution logic for known tasks.
@@
 (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic.
-Recent studies primarily aimed at improving logical accuracy by introducing more external validation such as self-consistency Narang et al.
+Recent studies primarily aimed at improving logical accuracy by external validation mechanisms like symbolic reasoning Cai et al.
+(2023); Pan et al.
+, stepwise verification including self-consistency Narang et al.
 ; Yu et al.
-(2024) or more hierarchical information such as Least-to-Most Zhou et al.
+(2024); Wang et al.
+(2024), self-refine Madaan et al.
+(2024), self-reflection Renze and Guven (2024) and more hierarchical information such as Least-to-Most Zhou et al.
 , Cumulative-Reasoning Zhang et al.
-and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors or poor flexibility.
+and Multi-experts Suzgun and Kalai (2024) strategies, but faced challenges related to cumulative errors Bao et al.
+(2024) or poor flexibility Yang et al.
+(2024a).
 Additionally, numerous studies also proposed more standardized recursive or backtracking branch forms from the logical data structure, including Tree-of-Thought (ToT) Yao et al.
@@
 (2024a).
-However, an efficient logical reasoning method that strikes a balance among logical accuracy, flexibility, and cost has yet to be discovered.
+However, an efficient logical reasoning method that strikes a balance among reasoning accuracy, flexibility, and cost has yet to be discovered.
 Our method is activated through meta cognition Fleur et al.
@@
 Integrating knowledge boundary within LLMs has emerged as a prospective strategy for enhancing their ability to avoid reasoning hallucinations of unknown knowledge through knowledge boundary constraints which requires additional algorithmic efforts Yin et al.
-(2024), external graph knowledge Tian et al.
+(2024); Chen et al.
+, external graph knowledge Tian et al.
 (2024), and training consumption Sun et al.
@@
 (2024) and reinforcement learning from human feedback (RLHF) Ouyang et al.
-(2022), instilling in LLMs a propensity for specific cognitive patterns, which manifests in two progressive layers of LLMs-taste description: (1) Stylistic template: encompassing grammatical and syntactic structures in descriptions.
+(2022), instilling in LLMs a propensity for specific cognitive patterns, which manifests in two progressive layers of LLMs-taste description: (1) Stylistic template: encompassing grammatical and syntactic structures in descriptions for thinking problems.
 (2) Solution logic: comprising problem-solving reasoning and methodological cues.
@@
 Inspired by RLHF Ouyang et al.
-(2022) utilizing preference data, and to derive high-cognitive preference prompt $P^{*}$ that enhance logical reasoning in LLMs, we propose a Preference-Guided Reverse Reasoning (PGRR) framework (detailed in alogrithm 1) mapping input-output demonstrations $D$ from an initial prompt $P$ to an optimal LLM-taste prompt $P_{opt}$.
+(2022) utilizing preference data, and to derive high-cognitive preference prompt $P^{*}$ that enhance logical reasoning in LLMs, we propose a Preference-Guided Reverse Reasoning (PGRR) framework (detailed in alogrithm 1) mapping input-output demonstrations $D$ from an initial prompt $P$ to an optimal LLM-taste prompt $P_{opt}$ from prompt candidates.
 (1) Reverse Reasoning Warm-up.
-We query the LLM $\mathcal{M}_{\text{LLM}}$ with a reversal prompt and demonstrations $\{P_{r},D\}$ (detailed in figure 3) for $warm$ iterations, generating a set of candidate cognitive preference responses $R=\{R_{1},R_{2},\dots,R_{warm}\}$ and their corresponding average probabilities $P_{res}^{i}$:
+We query the LLM $\mathcal{M}_{\text{LLM}}$ with a reversal prompt and demonstrations $\{P_{r},D\}$ (detailed in figure 4) for $warm$ iterations, generating a set of prompt candidates for solution $R=\{R_{1},R_{2},\dots,R_{warm}\}$ and their corresponding average probabilities $P_{res}^{i}$:
 $R$ $=\bigcup_{i=1}^{\text{warm}}R^{(i)}=\bigcup_{i=1}^{\text{warm}}\mathcal{M_{LLM}}(P_{r},D,i),$ (1)
@@
 $P_{pre}(R_{i+1}\succ R_{i})=\exp(\mathcal{M_{LLM}}(P_{eval},R_{i+1},R_{i}))$ (3)
-where $\mathcal{M_{LLM}}(Prompt_{eval},R_{i+1},R_{i})$ represents that require LLM to select more preferred data through $P_{eval}$ with a structure as "Please choose your more preferred instruction (A/B): (A).$R_{i+1}$; (B).$R_{i}$".
+where $\mathcal{M_{LLM}}(P_{eval},R_{i+1},R_{i})$ represents that require LLM to select more preferred data through $P_{eval}$ with a structure as "Please choose your more preferred instruction (A/B): (A).$R_{i+1}$; (B).$R_{i}$".
 Following the principle of preference transitivity Liu et al.
-(2024), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
+(2024b), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
 $P_{pre}(R_{i}\succ R_{j})=\begin{cases}1&i=j\\ \prod_{k=j}^{i-1}P_{pre}(R_{k+1}\succ R_{k})&i>j\\ 1-{P_{pre}(R_{j}\succ R_{i})}&i<j\end{cases}$ (4)
@@
 Within reverse reasoning, we further follow meta-cognitive Suzgun and Kalai (2024) using plan-and-solve by integrating logical algorithm pseudo-code to improve reasoning comprehension.
-And we incorporate fundamental mathematical logic symbols, including logical operators, quantifiers, inequalities and conditional statements, to facilitate model reasoning detailed in Figure 3.
+And we incorporate fundamental mathematical logic symbols, including logical operators, quantifiers, inequalities and conditional statements, to facilitate model reasoning detailed in Appendix A.2.
+3.3 Cognitive Preference Manager
+Algorithm 2 Cognitive Preference Manager (CPM)
+0:  $P$: Original prompt, $P^{*}$: Reverse-reasoned prompt, $\mathcal{M}_{emb}$: Offline LLM embedding model, $\delta$: Similarity threshold
+1:  $P_{task}\leftarrow P,P^{*}_{task}\leftarrow P^{*}$
+2:  $s\leftarrow\mathcal{M}_{emb}(P_{task},P^{*}_{task})$ Calculate embedding similarity between $P_{\text{task}}$ and $P^{*}_{\text{task}}$.
+3:  if $s\geq\delta$ then
+4:     Known detected: Enhance and refine $P$ to optimized instructions $P_{final}$.
+5:     Aggregate relevant task-specific knowledge.
+6:     return  $P_{final}$.
+7:  else
+8:     UnKnown detected: Adapt and expand $P_{task}$.
+9:     Leverage cognitive preference templates $T$ and $P$ to generate optimized instructions $P_{final}$.
+10:     return  $P_{final}$.
+11:  end if
+Cognitive Preference Manager.
+After reverse reasoning for LLMs-cognitive description $P^{*}$, We introduce an offline-deployed LLM embedding model $\mathcal{M}_{emb}$ to assist Cognitive Preference Manager (CPM) in determining whether reverse reasoning under reverse prompt $P_{r}$ and demonstrations $D$ reaches the knowledge boundary or cognitive error by calculating the similarity and setting a threshold $\delta$ (0.6$\sim$0.8 is recommended for optimal performance in distinguishing knowledge boundaries) between orginal task defination $P_{task}$ from $P$ and LLMs-cognitive task defination $P^{*}_{task}$ from $P^{*}$, and finally get a cognitive signal $S_{cog}$, formally:
+$S_{\text{cog}}=\begin{cases}unknown&,\texttt{sim}\left(\mathcal{M}_{\text{emb}}(P_{\text{task}}),\mathcal{M}_{\text{emb}}(P^{*}_{\text{task}})\right)<\delta\\ known&,\texttt{sim}\left(\mathcal{M}_{\text{emb}}(P_{\text{task}}),\mathcal{M}_{\text{emb}}(P^{*}_{\text{task}})\right)\geq\delta\end{cases}$ (7)
+where $\texttt{sim}(*)$ is a cosine similarity function that computes the similarity between two embedding vectors.
+By efficiently evaluating cognitive results, CPM integrates alternative aggregation strategies for $\mathcal{M_{LLM}}$ based on $S_{cog}$ 111we use $S_{cog}$ to conduct reverse evaluation of GPT-4 in our experiments., as detailed in appendix A.3: (1) Solution logic aggregation for known tasks: $\mathcal{M_{LLM}}$ merges beneficial aspects from the original prompt $P$ with the LLM-taste prompt $P^{*}$ to create the final prompt $P_{\textit{final}}$.
+(2) Stylistic template aggregation for unknown $\mathcal{M_{LLM}}$ extracts a cognitive preference template for thinking $T$ from the incorrect context in the LLM-taste prompt, and integrates meta-cognitive elements from the original prompt $P$ into $T$ to construct the final prompt $P_{final}$.
+Consequently, we utilize the final LLM-preferred prompt to query the LLM $\mathcal{M_{LLM}}$ with a specific problem input, $problem$, to obtain the final logical answer, $answer$:
+$answer=\mathcal{M_{LLM}}(P_{final},problem)$ (8)
+4 Experiments
+4.1 Datasets and Tasks
+Following Yao et al.
+(2024); Suzgun and Kalai (2024); Yang et al.
+(2024a), we evaluate RoT across eight tasks (detailed in Appendix A.1): (1) Mathematical Reasoning:Game of 24 Yao et al.
+(2024); Xiang et al.
+(2025), Multilingual Grade School Math Shi et al.
+(2022); (2) Domain Knowledge: Python Puzzles Schuster et al.
+(2021), Geometric Shapes, Multi-Step Arithmetic Two, Word Sorting, and Checkmate-in-One Srivastava et al.
+(2023); (3) Creativity: Shakespearean Sonnet Writing Suzgun and Kalai (2024).
+4.2 Baselines
+In our experiments, we compare RoT with five classic and latest state-of-the-art prompting baselines:
+  • CoT Prompting: Following Suzgun and Kalai (2024); Yang et al. (2024a), we employ GPT-4 to decompose instruction into logic intermediate reasoning steps activated by "Let’s think step by step".
+  • Meta-Prompting: Suzgun and Kalai (2024) introduced general, task-agnostic prompts as a scaffold to guide LLMs effectively perform logic tasks.
+  • Graph-of-Thought (GoT): Besta et al. (2024) modeled reasoning as a graph, where nodes are thoughts and edges define their relationships to solve complex problems.
+  • Tree-of-Thought (ToT): Yao et al.
+(2024) organized reasoning in a tree structure, allowing LLMs to explore multiple thought paths and select the most promising ones for problem-solving, enhancing their complexity management in logical reasoning tasks.
+  • Buffer of Thought (BoT): Yang et al.
+(2024a) introduced a meta-buffer that stores high-level thought templates, allowing for the adaptive retrieval and instantiation of relevant templates to enhance reasoning efficiency.
+4.3 Experiment Setup
+To ensure fair comparisons with previous methods, following Liu et al.
+(2024a); Yao et al.
+(2024); Zhang et al.
+; Suzgun and Kalai (2024), we utilize two LLMs: GPT-3.5-turbo and GPT-4, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study.
+For the $warmup$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API.
+Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency.
+For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^2] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al.
+(2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
+4.4 Evaluation Metrics
+Baselines Game of 24 Geometric Shapes Multi-Step Arithmetic Word Sorting Checkmate-in-One Python Puzzles MGSM Sonnet Writing
+GPT-3.5-turbo
+CoT (NeurIPS) Wei et al. (2022) 1-shot 26.3 50.2 73.9 71.9 25.7 30.1 75.0 58.6
+2-shot 29.6 53.2 76.2 74.6 25.4 30.6 77.4 61.3
+Meta-prompting Suzgun and Kalai (2024) 1-shot 59.6 69.5 83.1 90.0 47.4 40.2 80.3 69.5
+2-shot 60.2 71.0 85.1 91.5 48.0 41.1 81.5 71.4
+ToT (NeurIPS) Yao et al. (2024) 1-shot 66.7 56.2 81.9 89.2 44.2 38.2 78.3 63.5
+2-shot 67.3 58.5 82.3 90.3 45.2 39.5 79.4 65.3
+GoT (AAAI) Besta et al. (2024) 1-shot 69.7 58.6 79.1 88.8 45.4 35.3 80.3 62.2
+2-shot 70.4 58.9 79.0 89.5 46.0 36.7 81.0 63.7
+BoT (NeurIPS) Yang et al. (2024a) 1-shot 74.7 83.1 87.6 93.2 67.9 48.2 81.1 71.1
+2-shot 75.5 84.3 88.3 94.4 68.5 48.0 82.3 73.0
+RoT (Ours) 1-shot 82.8 88.4 89.2 95.2 71.5 50.2 82.3 75.5
+2-shot 87.8 88.7 89.5 95.6 72.6 50.4 84.7 75.8
+GPT-4
+CoT (NeurIPS) Wei et al. (2022) 1-shot 31.3 57.4 83.1 80.7 35.8 35.3 83.9 66.3
+2-shot 32.7 60.9 85.5 82.7 36.7 35.9 84.3 70.2
+Meta-prompting Suzgun and Kalai (2024) 1-shot 64.6 76.7 89.6 97.6 58.2 45.0 85.1 78.7
+2-shot 66.3 78.2 90.7 98.4 58.9 45.6 85.5 79.8
+ToT (NeurIPS) Yao et al. (2024) 1-shot 73.7 58.6 88.8 95.2 49.8 43.8 84.7 68.7
+2-shot 74.5 60.1 90.3 96.0 48.8 45.2 86.3 69.8
+GoT (AAAI) Besta et al. (2024) 1-shot 74.7 56.2 87.6 95.6 49.0 42.6 85.9 68.3
+2-shot 73.5 57.7 88.7 96.8 51.2 43.5 86.7 62.5
+BoT (NeurIPS) Yang et al. (2024a) 1-shot 82.8 90.4 94.8 99.2 87.1 51.8 87.9 79.1
+2-shot 83.7 90.7 96.8 99.6 88.3 52.8 87.5 79.8
+RoT (Ours) 1-shot 97.0 94.8 98.4 99.6 91.2 54.6 88.7 89.2
+2-shot 98.0 95.2 99.2 100.0 92.0 57.3 90.0 92.0
+Table 1: Comparison of RoT with baselines across 1-shot and 2-shot settings for reasoning accuracy.
+Knowledge Boundary and Cognitive Preference Consistency.
+As described in Section 3.3, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$.
+We first evaluate GPT-3.5-turbo & GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^3] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot.
+and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
+Reasoning Accuracy.
+Following Suzgun and Kalai (2024); Yao et al.
+(2024), we introduce a LLM444we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong).
+We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
+Reasoning Efficiency.
+Following Yang et al. (2024a), we evaluate reasoning efficiency in terms of complexity by calculating $T$, calculated as the average time spent per task across all samples in the test dataset:
+$T=\frac{1}{N}\sum_{i=1}^{M}\sum_{j=1}^{N}T_{ij}$ (9)
+where $N$ is the total number of tasks, $M$ is the total number of samples, and $T_{ij}$ represents the time taken for the $i$-th task on the $j$-th sample.
+5 Results and Discussion
+5.1 Knowledge Boundary and Cognitive Preference Consistency
+Task 1-shot 2-shot
+Kno. Unkno. Kno. Unkno.
+Game of 24
+MGSM (avg)
+Multi-Step Arithmetic
+WordSorting
+Python Puzzles
+Geometric Shapes
+Checkmate-in-One
+Sonnet Writing
+Table 2: Knowledge Boundary under Reversal Reasoning for GPT-3.5-turbo and GPT-4 with identical results. Kno. for Known, Unkno. for Unknown.
+Knowledge Boundary.
+As shown in Table 2, our findings provide preliminary support for our hypothesis concerning cognitive knowledge boundaries of LLMs in reverse reasoning across current task benchmarks in one-shot and two-shot settings Yao et al.
+(2024); Suzgun et al.
+(2023); Srivastava et al.
+(2023); Schuster et al.
+(2021); Shi et al.
+(2022); Suzgun and Kalai (2024).
+Notably, GPT-3.5-turbo & GPT-4 excel in structured reasoning tasks such as Game of 24, Geometric Shapes, and Checkmate-in-One, consistently achieving strong results in both one-shot and two-shot settings.
+In contrast, tasks such as MGSM (avg) and Python Puzzles fall into an unknown category in the one-shot scenario, stemming from multi-source problems that are challenging to make reverse reasoning without sufficient context, indirectly supporting the rationale for our CPM approach.
+Cognitive Preference Consistency.
+Our human evaluation of $\textit{Con}_{cog}$ yielded high scores of 4.32 in the 2-shot setting and 4.01 in the 1-shot setting, further validating that LLMs exhibit cognitive preferences shaped by RLHF Ouyang et al.
+(2022).
+5.2 Reasoning Accuracy and Efficiency
+(1) RoT can activate LLMs’ reasoning accuracy.
+As shown in Table 1, RoT consistently outperforms all baselines across various reasoning tasks, with particularly notable improvements observed in GPT-4.
+Specifically, compared with the best BoT, RoT achieves significant gains in tasks such as Game of 24 (+17.15% in 1-shot and +17.08% in 2-shot), Geometric Shapes (+4.87% in 1-shot and +4.96% in 2-shot), and Checkmate-in-One (+4.71% in 1-shot and +4.19% in 2-shot), demonstrating our flexibility of leveraging cognitive preference in LLM to activate logic capabilities through reverse reasoning (case study detailed in appendix B).
+For GPT-3.5-turbo, RoT also demonstrates substantial improvements, such as Game of 24 (+10.84% in 1-shot and +16.29% in 2-shot) and Geometric Shapes (+6.38% in 1-shot and +5.22% in 2-shot), further emphasizing its versatility in activating logic through cognitive preference management.
+Figure 3: Inference time comparison
+Ablation study Game of 24 Geometric Shapes Multi-Step Arithmetic Word Sorting Checkmate-in-One Python Puzzles×$\times$ MGSM×$\times$ Sonnet Writing
+GPT-3.5-turbo
+RoT 1-shot 82.8 88.4 89.2 95.2 71.5 50.2 82.3 75.5
+2-shot 87.8 88.7 89.5 95.6 72.6 50.4 84.7 75.8
+w/o PGRR 1-shot 74.7 72.3 74.3 74.7 65.5 44.2 78.7 69.6
+2-shot 75.5 72.6 75.4 76.2 66.9 44.4 79.0 69.8
+w/o Logic 1-shot 76.7 80.7 81.5 84.3 66.3 45.8 77.5 72.9
+2-shot 76.5 80.2 81.9 85.1 66.5 46.0 78.2 73.0
+w/o CPM 1-shot 80.8 86.0 88.4 93.2 70.3 39.8 72.7 74.2
+2-shot 81.6 86.7 88.3 93.5 71.0 47.2 80.2 75.0
+GPT-4
+RoT 1-shot 97.0 94.8 98.4 99.6 91.2 54.6 88.7 89.2
+2-shot 98.0 95.2 99.2 100.0 92.0 57.3 90.0 92.0
+w/o PGRR 1-shot 77.8 74.7 78.7 79.5 80.7 45.0 85.9 72.7
+2-shot 78.6 75.0 79.0 81.0 82.3 46.8 86.0 73.0
+w/o Logic 1-shot 85.9 83.1 87.1 89.2 81.5 49.8 85.5 86.1
+2-shot 86.7 83.5 87.9 89.5 83.5 50.0 86.7 77.0
+w/o CPM 1-shot 92.0 90.0 94.8 97.6 89.2 40.6 84.7 88.8
+2-shot 93.9 90.3 95.2 98.4 89.1 51.6 86.3 91.5
+Table 2: Ablation study of RoT across various tasks in 1-shot and 2-shot settings. indicates known tasks in both 1-shot and 2-shot; $\times$ indicates unknown tasks in 1-shot.
+(2) RoT demonstrates better tradeoff between reasoning accuracy and efficiency across baselines.
+As shown in Figure 3, RoT achieves competitive performance in reasoning efficiency, outperforming baselines and being second only to BoT, while BoT’s dependence on numerous pre-defined golden thought templates limits its flexibility.
+In contrast, our RoT, as a plugin strategy, emphasizes exploring optimal prompt for solution and task data instantiation after reverse reasoning warm-up.
+5.3 Ablation Study
+As shown in Table 2, we conducted three ablation studies to evaluate key components: (1) w/o PGRR: Removing Preference-Guided Reverse Reasoning(PGRR) for exploring LLMs-taste prompts; (2) w/o Logic: Excluding mathematical logic for pseudo-code plan-and-solve; and (3) w/o CPM: Eliminating Cognitive Preference Manage(CPM) for both known and unknown tasks.
+Impact of PGRR.
+Excluding Preference-Guided Reverse Reasoning (PGRR) results in a significant reduction in overall task performance, as seen in tasks like the Game of 24 (98.0% to 78.6% for GPT-4, 87.8% to 75.5% for GPT-3.5-turbo) and WordSorting (100% to 81% for GPT-4, 95.6% to 76.2% for GPT-3.5-turbo), indicating w/o PGRR weaken model’s cognition to task-specific requirements.
+Impact of Logic.
+Removing mathematical logic from RoT leads to notable declines in tasks requiring structured problem-solving, such as Multi-Step Arithmetic (99.2% to 87.9% for GPT-4, 89.5% to 81.9% for GPT-3.5-turbo) and Checkmate-in-One (92.0% to 83.5% for GPT-4, 72.6% to 66.5% for GPT-3.5-turbo), underscoring w/o Logic negatively affects structured problem-solving for complex reasoning.
+Impact of CPM.
+Lacking Cognitive Preference Manager (CPM) has a particularly pronounced effect on unknown tasks, with Python Puzzles dropping from 54.6% to 45.0% for GPT-4 and 50.2% to 39.8% for GPT-3.5-turbo, while MGSM decreased from 84.7% to 80.2% for GPT-3.5-turbo and from 90.0% to 86.3% for GPT-4, indicating w/o CPM weaken RoT’s flexibility for tackling known tasks and unknown tasks  [^5] Please refer to Appendix B.1.1 for case study on known task and Appendix B.2.1 for case study on unknwon task..
+6 Conclusion
+In this paper, we propose Reversal of Thought (RoT), a novel and plug-and-play framework to enhance the logical reasoning capabilities of LLMs.
+By integrating reverse reasoning with meta-cognitive mechanisms and cognitive preference management, RoT improves reasoning accuracy and efficiency while minimizing computational costs, which leverages Preference-Guided Reverse Reasoning and Cognitive Preference Manager, which optimally aligns LLM reasoning processes with their cognitive preferences shaped by their pretraining and RLHF.
+Comprehensive experiments across diverse reasoning tasks demonstrate that RoT consistently outperforms state-of-the-art baselines in both known and unknown task scenarios, demonstrating the potential to expand knowledge boundaries through cognitive preference template.
+Our research provides valuable insights into future studies focused on further enhancing LLMs’ reasoning capacities by dynamically exploring cognitive preferences for complex reasoning tasks.
+Limitations
+Reversal of Thought (RoT) introduce a reverse reasoning warm-up to activate cognitive preference for LLMs to enhance logic capabilities and introduce a cognitive preference manager to determine knowledge boundary and utilize cognitive preference for known and unknown tasks.
+While RoT has performed exceptionally in logic accuracy and efficiency, We discuss major challenge in its reliance on two-shot demonstration inputs involving two distinct problem cases.
+We observed that RoT may struggles with one-shot learning in multi-source tasks.
+we partially and effectively mitigates this issue through the integration of Cognitive Preference Manager (CPM) and two-shot learning.
+In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks.
+Furthermore, we believe that utilizing Reversal of Thought in teacher-student model distillation, which could further amplify the practical value of our approach.
+Ethical Considerations
+Since the datasets used in our experiments focus on pure mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity, which are all sourced from publicly available datasets Yao et al.
+(2024); Suzgun et al.
+(2023); Srivastava et al.
+(2023); Schuster et al.
+(2021); Shi et al.
+(2022); Suzgun and Kalai (2024); Xiang et al.
+(2025) and devoid of any personal privacy or sensitive ethical information.
+Therefore, we do not identify any immediate ethical concerns regarding our current work.
+Additionally, we conduct human evaluations of cognitive preference consistency with three anonymous professional annotators following our instructions (detailed in appendix C).
+Appendix A Experimental Details
+A.1 Datasets and Tasks
+To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al.
+(2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al.
+(2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al.
+(2024); Xiang et al.
+(2025) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24666https://www.4nums.com/game/ .
+(2) BIG-Bench Suzgun et al.
+(2023); Srivastava et al.
+(2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting and Checkmate-in-One from BIG-Bench suite  [^7] https://huggingface.co/datasets/google/bigbench Srivastava et al.
+(2023); (3) Python Puzzles Schuster et al.
+(2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels [^8] https://github.com/microsoft/PythonProgrammingPuzzles;(4) Multilingual Grade School Math (MGSM) Shi et al.
+(2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al.
+(2021), featuring translations of a subset of examples into ten diverse languages [^9] https://github.com/google-research/url-nlp/tree/main/mgsm .(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three  [^10] https://huggingface.co/datasets/turingmachine/meta-prompting or five  [^11] https://github.com/iljones00/Shakespearean-Sonnets-GPT specified words verbatim.
+A.2 Prompt for Reverse Reasoning
 Prompt for Reverse Reasoning Instruction
@@
 Figure 4: Prompt for Reverse Reasoning
-3.3 Cognitive Preference Manager
-Algorithm 2 Cognitive Preference Manager (CPM)
-0:  $P$: Original prompt, $P^{*}$: Reverse-reasoned prompt, $\mathcal{M}_{emb}$: Offline LLM embedding model, $\delta$: Similarity threshold
-1:  $P_{task}\leftarrow P,P^{*}_{task}\leftarrow P^{*}$
-2:  $s\leftarrow\mathcal{M}_{emb}(P_{task},P^{*}_{task})$ Calculate embedding similarity between $P_{\text{task}}$ and $P^{*}_{\text{task}}$.
-3:  if $s\geq\delta$ then
-4:     Known detected: Enhance and refine $P$ to optimized instructions $P_{final}$.
-5:     Aggregate relevant task-specific knowledge.
-6:     return  $P_{final}$.
-7:  else
-8:     Unknown detected: Adapt and expand $P_{task}$.
-9:     Leverage cognitive preference templates $T$ and $P$ to generate optimized instructions $P_{final}$.
-10:     return  $P_{final}$.
-11:  end if
-Cognitive Preference Manager.
-After reverse reasoning for LLMs-cognitive description $P^{*}$, We introduce an offline-deployed LLM embedding model $\mathcal{M}_{emb}$ to assist Cognitive Preference Manager (CPM) in determining whether reverse reasoning under reverse prompt $P_{r}$ and demonstrations $D$ reaches the knowledge boundary or cognitive error by calculating the similarity and setting a threshold $\delta$ (0.6$\sim$0.8 is recommended for optimal performance in distinguishing knowledge boundaries) between orginal task defination $P_{task}$ from $P$ and LLMs-cognitive task defination $P^{*}_{task}$ from $P^{*}$, and finally get a cognitive signal $S_{cog}$, formally:
-$S_{\text{cog}}=\begin{cases}unknown&,\texttt{sim}\left(\mathcal{M}_{\text{emb}}(P_{\text{task}}),\mathcal{M}_{\text{emb}}(P^{*}_{\text{task}})\right)<\delta\\ known&,\texttt{sim}\left(\mathcal{M}_{\text{emb}}(P_{\text{task}}),\mathcal{M}_{\text{emb}}(P^{*}_{\text{task}})\right)\geq\delta\end{cases}$ (7)
-where $\texttt{sim}(*)$ is a cosine similarity function that computes the similarity between two embedding vectors.
-By efficiently evaluating cognitive results, CPM integrates alternative aggregation strategies for $\mathcal{M_{LLM}}$ based on $S_{cog}$ 111we use $S_{cog}$ to conduct reverse evaluation of GPT-4 in our experiments., as detailed in appendix A: (1) Solution logic aggregation for known tasks: $\mathcal{M_{LLM}}$ merges beneficial aspects from the original prompt $P$ with the LLM-taste prompt $P^{*}$ to create the final prompt $P_{\textit{final}}$.
-(2) Stylistic template aggregation for unknown $\mathcal{M_{LLM}}$ extracts a cognitive preference template $T$ from the incorrect context in the LLM-taste prompt, and integrates meta-cognitive elements from the original prompt $P$ into $T$ to construct the final prompt $P_{final}$.
-Consequently, we utilize the final LLM-preferred prompt to query the LLM $\mathcal{M_{LLM}}$ with a specific problem input, $problem$, to obtain the final logical answer, $answer$:
-$answer=\mathcal{M_{LLM}}(P_{final},problem)$ (8)
-4 Experiments
-4.1 Datasets and Tasks
-To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al.
-(2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al.
-(2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al.
-(2024) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24.
-(2) BIG-Bench Suzgun et al.
-(2023); Srivastava et al.
-(2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting from BIG-Bench Hard Suzgun et al.
-(2023) and Checkmate-in-One from BIG-Bench suite Srivastava et al.
-(2023); (3) Python Puzzles Schuster et al.
-(2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels.;(4) Multilingual Grade School Math (MGSM) Shi et al.
-(2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al.
-(2021), featuring translations of a subset of examples into ten diverse languages.(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three specified words verbatim.
-4.2 Baselines
-In our experiments, we compare RoT with five classic and latest state-of-the-art prompting baselines:
-  • CoT Prompting: Following Suzgun and Kalai (2024); Yang et al. (2024a), we employ GPT-4 to decompose instruction into logic intermediate reasoning steps activated by "Let’s think step by step".
-  • Meta-Prompting: Suzgun and Kalai (2024) introduced general, task-agnostic prompts as a scaffold to guide LLMs effectively perform logic tasks.
-  • Graph-of-Thought (GoT): Besta et al. (2024) modeled reasoning as a graph, where nodes are thoughts and edges define their relationships to solve complex problems.
-  • Tree-of-Thought (ToT): Yao et al.
-(2024) organized reasoning in a tree structure, allowing LLMs to explore multiple thought paths and select the most promising ones for problem-solving, enhancing their complexity management in logical reasoning tasks.
-  • Buffer of Thought (BoT): Yang et al.
-(2024a) introduced a meta-buffer that stores high-level thought templates, allowing for the adaptive retrieval and instantiation of relevant templates to enhance reasoning efficiency.
-4.3 Experiment Setup
-To ensure fair comparisons with previous methods, we utilize GPT-4, the latest model supporting $logprobs$  [^2] https://platform.openai.com/docs/api-reference/introduction, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study.
-For the $warm$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API.
-Our findings suggest that a value of 5 optimally balances between logical accuracy and cost-efficiency.
-For embedding model $\mathcal{M}_{emb}$, we utilize a huggingface model stella_en_1.5B_v5  [^3] https://huggingface.co/dunzhang/stella_en_1.5B_v5, License:mit, a high-performance model with the smallest parameter countamong the top three on the MTEB leaderboard Muennighoff et al.
-(2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
-4.4 Evaluation Metrics
-Task CoT (NeurIPS) Wei et al. (2022) Meta-prompting Suzgun and Kalai (2024) ToT (NeurIPS) Yao et al. (2024) GoT (AAAI) Besta et al. (2024) BoT (NeurIPS) Yang et al. (2024a) RoT (Ours)
-1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot
-Game of 24 31.5 33.2 65.4 66.8 73.3 74.1 73.0 73.5 82.8 82.5 97.5 97.8
-Geometric Shapes 57.3 60.8 76.8 78.3 58.8 60.1 56.1 57.8 90.1 90.9 94.8 95.0
-Multi-Step Arithmetic 83.4 85.3 89.5 90.6 88.5 90.3 87.5 88.8 94.6 96.8 98.6 99.0
-WordSorting 80.8 82.5 97.8 98.6 95.0 95.8 95.6 96.8 99.2 99.6 99.8 100.0
-Checkmate-in-One 35.8 36.5 58.0 57.5 49.8 48.8 49.0 51.2 87.2 88.1 91.8 92.0
-Python Puzzles 35.3 35.8 45.0 45.5 43.8 45.1 42.5 43.5 51.8 52.6 54.8 57.5
-MGSM (avg) 83.8 84.2 85.2 85.4 84.5 86.1 85.8 86.8 87.5 87.3 88.9 90.1
-Sonnet Writing 66.5 70.2 78.8 79.8 68.9 69.8 68.0 62.5 79.0 79.5 89.5 92.0
-Table 1: Comparison of RoT with baselines across 1-shot and 2-shot settings for reasoning accuracy.
-Knowledge Boundary and Cognitive Preference Consistency.
-As described in Section 3.2, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$.
-We first evaluate GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^4] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot.
-and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
-Reasoning Accuracy.
-Following Suzgun and Kalai (2024); Yao et al.
-(2024), we introduce a LLM555we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong).
-We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
-Reasoning Efficiency.
-We evaluate reasoning efficiency by calculating ${t_{eff}}$, calculated as the average time spent per task across all samples in the test dataset:
-$t_{eff}=\frac{1}{N}\sum_{j=1}^{M}\sum_{i=1}^{N}t_{ij}$ (9)
-where $N$ is the total number of tasks, $M$ is the total number of samples, and $t_{ij}$ represents the time taken for the $i$-th task on the $j$-th sample.
-5 Results and Discussion
-Task 1-shot 2-shot
-Kno. Unkno. Kno. Unkno.
-Game of 24
-MGSM (avg)
-Multi-Step Arithmetic
-WordSorting
-Python Puzzles
-Geometric Shapes
-Checkmate-in-One
-Sonnet Writing
-Table 2: Knowledge Boundary under Reversal Reasoning. Kno. for Known, Unkno. for Unknown.
-5.1 Knowledge Boundary and Cognitive Preference Consistency
-Knowledge Boundary.
-As shown in Table 2, our findings provide preliminary support for our hypothesis concerning cognitive knowledge boundaries of LLMs in reverse reasoning across current task benchmarks in one-shot and two-shot settings Yao et al.
-(2024); Suzgun et al.
-(2023); Srivastava et al.
-(2023); Schuster et al.
-(2021); Shi et al.
-(2022); Suzgun and Kalai (2024).
-Notably, GPT-4 excels in structured reasoning tasks such as Game of 24, Geometric Shapes, and Checkmate-in-One, consistently achieving strong results in both one-shot and two-shot settings.
-In contrast, tasks such as MGSM (avg) and Python Puzzles fall into an unknown category in the one-shot scenario, stemming from multi-source problems that are challenging to make reverse reasoning without sufficient context, indirectly supporting the rationale for our CPM approach, detailed in appendix B.
-Cognitive Preference Consistency.
-Moreover, our human evaluation of $\textit{Con}_{cog}$ produced a high score of 4.32 in the two-shot setting and 4.01 in the one-shot setting, further validating that LLMs exhibit cognitive preferences shaped by RLHF Ouyang et al.
-(2022).
-5.2 Reasoning Accuracy and Efficiency
-Reasoning Accuracy.
-As shown table 1, our RoT model consistently outperforms all baseline models across various logic reasoning tasks.
-Notably, compared with the best BoT, we observe significant accuracy improvements in 1-shot and 2-shot settings, particularly evidenced by increases of 14.7% and 15.3% on the Game of 24, 4.7% and 4.1% on Geometric Shapes, and 4.6% and 4.5% on Checkmate-in-One, demonstrating our flexibility and efficiency of leveraging cognitive preference in LLMs to activate logic capabilities through reverse reasoning, particularly with cognitive preference management.
-Reasoning Efficiency.
-Figure 4: Inference time comparison
-Dataset RoT w/o PGRR w/o Logic w/o CPM
-1-shot 2-shot 1-shot 2-shot 1-shot 2-shot 1-shot 2-shot
-Game of 24 97.5 97.8 77.5 77.8 86.0 86.5 92.0 93.5
-Geometric Shapes 94.8 95.0 74.8 75.0 83.0 83.5 90.0 90.5
-Multi-Step Arithmetic 98.6 99.0 78.6 79.0 87.5 88.0 95.0 95.5
-WordSorting 99.8 100.0 79.1 81.0 89.0 89.2 98.0 98.5
-Checkmate-in-One 91.8 92.0 80.8 82.0 81.0 83.5 98.0 88.5
-$\times$×Python Puzzles 54.8 57.5 44.3 47.1 50.0 50.5 41.8 52.5
-$\times$×MGSM (avg) 88.9 90.1 85.8 86.1 86.0 86.5 85.2 86.1
-Sonnet Writing 89.5 92.0 72.5 73.0 87.0 77.5 89.0 91.5
-Table 3: Ablation study of RoT across various tasks in 1-shot and 2-shot settings. indicates known tasks in both 1-shot and 2-shot; $\times$ indicates unknown tasks in 1-shot.
-As shown in Figure 4, RoT achieves competitive performance in reasoning efficiency, outperforming baselines and being second only to BoT, while BoT requires numerous initial golden thought templates in advance, which subtly limits its flexibility.
-In contrast, RoT focuses on exploring LLM-taste prompts, requiring only data for reasoning in subsequent steps after reverse reasoning warm-up.
-5.3 Ablation Study
-As shown in Table 3, we conducted three ablation studies to evaluate key components: (1) w/o PGRR: Removing Preference-Guided Reverse Reasoning(PGRR) for exploring LLMs-taste prompts; (2) w/o Logic: Excluding mathematical logic for pseudo-code plan-and-solve; and (3) w/o CPM: Eliminating Cognitive Preference Manage(CPM) for both known and unknown tasks.
-Impact of PGRR.
-Excluding Preference-Guided Reverse Reasoning (PGRR) for exploring LLMs-taste prompts regarding tasks results in a significant reduction in overall task performance, with decreases of up to 20% in tasks like the Game of 24 (97.8% to 77.8%) and WordSorting (100% to 81%), indicating w/o PGRR weaken model’s cognition to task-specific requirements.
-Impact of Logic.
-Removing mathematical logic from RoT leads to notable declines in tasks requiring structured problem-solving, such as Multi-Step Arithmetic (99.0% to 88.0%) and Checkmate-in-One (92.0% to 83.5%), underscoring w/o Logic negatively affects structured problem-solving for complex reasoning.
-Impact of CPM.
-Lacking Cognitive Preference Manager (CPM) has a particularly pronounced effect on unknown tasks with Python Puzzles dropping from 57.5% to 52.5% and MGSM from 90.1% to 86.1%, indicating w/o CPM weaken RoT’s flexibility for tackling known tasks (case study for knwon task detailed in appendix B.1) and unknwon tasks (case study for unknwon task is detailed in Appendix B.2).
-6 Conclusion
-In this paper, we propose Reversal of Thought (RoT), a novel framework to enhance the logical reasoning capabilities of LLMs.
-By integrating reverse reasoning with meta-cognitive mechanisms and cognitive preference management, RoT improves reasoning accuracy and efficiency while minimizing computational costs, which leverages Preference-Guided Reverse Reasoning and Cognitive Preference Manager, which optimally aligns LLM reasoning processes with their cognitive preferences shaped by their pretraining and RLHF.
-Comprehensive experiments across diverse reasoning tasks demonstrate that RoT consistently outperforms state-of-the-art baselines in both known and unknown task scenarios, demonstrating the potential to expand knowledge boundaries through cognitive preference template.
-Our research provides valuable insights into future studies focused on further enhancing LLMs’ reasoning capacities by dynamically exploring cognitive preferences for complex reasoning tasks.
-Limitations
-Reversal of Thought (RoT) introduce a reverse reasoning warm-up to activate cognitive preference for LLMs to enhance logic capabilities and introduce a cognitive preference manager to determine knowledge boundary and utilize cognitive preference for known and unknown tasks.
-While RoT has performed exceptionally in logic accuracy and efficiency, We discuss major challenge in its reliance on two-shot demonstration inputs involving two distinct problem cases.
-We observed that RoT may struggles with one-shot learning in multi-source tasks.
-we partially and effectively mitigates this issue through the integration of Cognitive Preference Manager (CPM) and two-shot learning.
-In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks.
-Ethical Considerations
-Since the datasets used in our experiments focus on pure mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity, which are all sourced from publicly available datasets Yao et al.
-(2024); Suzgun et al.
-(2023); Srivastava et al.
-(2023); Schuster et al.
-(2021); Shi et al.
-(2022); Suzgun and Kalai (2024) and devoid of any personal privacy or sensitive ethical information.
-Therefore, we do not identify any immediate ethical concerns regarding our current work.
-Additionally, we conduct human evaluations of cognitive preference consistency with three anonymous professional annotators following our instructions (detailed in appendix C), and they receive equivalent compensation.
-Appendix A Prompt for CPM (Known/Unknown)
+A.3 Prompt for CPM (Known/Unknown)
 Prompt for CPM (Known/Unknown) Instruction
@@
 Output: [Output]
+Figure 5: Prompt for CPM (Known/Unknown)
 Appendix B Case Study
 B.1 Case Study for Known Task (Game of 24)
+B.1.1 $P_{final}$ Case study for Game of 24
 $P_{final}$ Case study for known task after CPM Task Defination
@@
 Logical Pseudocode
-$\forall P(A,B,C,D)$ in permutations of [A, B, C, D].
-$\forall op_{n}$ in (+, -, $\times$, $\div$) where $n\in\{0,1,2,3\}$.
-For each permute $\neg\exists(Expression==24)\lor(Expression=permute[0]\;op_{1}\;permute[1]\\ ;op_{2}\;permute[2]\;op_{3}\;permute[3])$: $\exists(Expression==24)\supset Print(Expression)\land Terminate\;Process.$ Print("No feasible solution exists.").
-Case Examples: Example 1: Input: 4 6 7 1 Output: One feasible solution (e.g., $6/(1-(4/7))=24$).
-Example 2: Input: 4 2 2 1 Output: "No feasible solution exists." Input-Output Format:
+[…]
 Input: Four integers separated by space (e.g., "3 3 8 8").
@@
 If no solution exists, output the string "No feasible solution exists."
+B.1.2 Instantiation Case Study for Game of 24 Using $P_{final}$
+Instantiation Case Study for Game of 24 Using $P_{final}$ Input: 4 9 10 13
+Steps: Let me compute possible expressions for this input that result in 24 using provided logic pseudocode.
+1. Generate all permutations of the numbers (4,9,10,13): This gives 24 possible arrangements.
+2. Apply operations: For each permutation, test all combinations of the operations (+, -, *, /) between the numbers.
+3. Check if the result equals 24: Evaluate the expression and check if it equals 24.
+4. Output the result: If a valid expression is found, print it. If not, output that no feasible solution exists.
+import itertools
+inputs = [4,9,10,13]
+operations = [’+’, ’-’, ’*’, ’/’]
+def evaluate_expression(expr):
+    try:
+        return eval(expr)
+    except ZeroDivisionError:
+        return None
+permutations = itertools.permutations(inputs)
+op_combinations = itertools.product(operations, repeat=3)
+solutions = []
+for perm in permutations:
+    for ops in op_combinations:
+        exprs = [
+            f"({perm[0]} {ops[0]} {perm[1]}) {ops[1]} ({perm[2]} {ops[2]} {perm[3]})",
+            f"({perm[0]} {ops[0]} ({perm[1]} {ops[1]} {perm[2]})) {ops[2]} {perm[3]}",
+            f"{perm[0]} {ops[0]} ({perm[1]} {ops[1]} ({perm[2]} {ops[2]} {perm[3]}))",
+            f"({perm[0]} {ops[0]} {perm[1]}) {ops[1]} ({perm[2]} {ops[2]} {perm[3]})",
+        for expr in exprs:
+            if evaluate_expression(expr) == 24:
+                solutions.append(expr)
+print(solutions)
+  Output: 4 * (9 + (10 - 13))=24
 B.2 Case Study for Unknown Task (MGSM)
+B.2.1 $P_{final}$ Case study for MGSM
 $P_{final}$ Case study for unknown task after CPM Task Defination
@@
 Output: A single integer or float number that is the result of the operations described in the scenario.
+B.2.2 Instantiation Case Study for MGSM Using $P_{final}$
+Instantiation Case Study for MGSM Using $P_{final}$ Input: Paul is at a train station and is waiting for his train.
+He isn’t sure how long he needs to wait, but he knows that the fourth train scheduled to arrive at the station is the one he needs to get on.
+The first train is scheduled to arrive in 10 minutes, and this train will stay in the station for 20 minutes.
+The second train is to arrive half an hour after the first train leaves the station, and this second train will stay in the station for a quarter of the amount of time that the first train stayed in the station.
+The third train is to arrive an hour after the second train leaves the station, and this third train is to leave the station immediately after it arrives.
+The fourth train will arrive 20 minutes after the third train leaves, and this is the train Paul will board.
+In total, how long, in minutes, will Paul wait for his train?
+Steps:
+Step 1: Symbol Information
+1. $\alpha$ (first train arrival) = 10 minutes
+2. $\beta$ (first train stay) = 20 minutes
+3. $\gamma$ (second train stay as a quarter of the first train’s stay) = $\frac{\beta}{4}$
+4. $\delta$ (time between second train departure and third train arrival) = 60 minutes
+5. $\varepsilon$ (fourth train arrival after third train leaves) = 20 minutes
+Step 2: Identify Contextual Clues and Operations
+1. First train departure time = $\alpha+\beta$
+2. Second train arrival time = First train departure time + 30 minutes
+3. Second train stay time = $\gamma$
+4. Third train arrival time = Second train departure time + $\delta$
+5. Fourth train arrival time = Third train departure time + $\varepsilon$
+Step 3: Perform Calculations Based on Identified Operations
+[…]
+2. Second train arrival time = Add(First train departure time,30) = $30+30=60$ minutes
+[…]
+4. Second train departure time = Add(Second train arrival time ,Second train stay time) = $60+5=65$ minutes
+5. Third train arrival time = Add(Second train departure time,$\delta$) = $65+60=125$ minutes
+6. Third train departure time = Third train arrival time (since it leaves immediately) = 125 minutes
+7. Fourth train arrival time = Add(Third train departure time, $\varepsilon$) = $125+20=145$ minutes
+Step 4: Calculate the Final Result
+Final Result: Fourth train arrival time = 145 minutes
+  Output: 145 minutes
 Appendix C Instructions for Human Evaluation.
@@
 Criteria.
-(1) Style Consistency: How well responses maintain a consistent stylistic tone across different tasks. (2) Coherence: Internal logic and smooth flow of the response.
+(1) Style Consistency: How well responses maintain a consistent stylistic tone for thinking across different tasks. (2) Coherence: Internal logic and smooth flow of the response.
 Scoring.

### v3

@@
 Reversal of Thought: Enhancing Large Language Models with Preference-Guided Reverse Reasoning Warm-up
-Jiahao Yuan1, Dehui Du1, Hao Zhang1, Zixiang Di1, Usman Naseem2 1East China Normal University 2Macquarie University 51275900024@stu.ecnu.edu.cn, dhdu@sei.ecnu.edu.cn, {71275902039,51265901113}@stu.ecnu.edu.cn, usman.naseem@mq.edu.au
+Jiahao Yuan1thanks: 51275900024@stu.ecnu.edu.cn, Dehui Du1thanks: Corresponding Author: dhdu@sei.ecnu.edu.cn, Hao Zhang1, Zixiang Di1, Usman Naseem2 1East China Normal University 2Macquarie University
 1 Introduction
@@
 ; Wang et al.
-(2024) and recursive reasoning process Lee and Kim (2023); Yu et al.
+(2024a) and recursive reasoning process Lee and Kim (2023); Yu et al.
 (2024) through Tree-of-Thought (ToT) Yao et al.
@@
 (2022) has been proven to be a promising approach that incorporates an intermediate logic chain to enhance LLMs’ logic.
-Recent studies primarily aimed at improving logical accuracy by external validation mechanisms like symbolic reasoning Cai et al.
+Recent studies primarily aimed at improving logical accuracy by external validation mechanisms Bi et al.
+(2025) like symbolic reasoning Cai et al.
 (2023); Pan et al.
@@
 (2024); Wang et al.
-(2024), self-refine Madaan et al.
+(2024a), self-refine Madaan et al.
 (2024), self-reflection Renze and Guven (2024) and more hierarchical information such as Least-to-Most Zhou et al.
@@
 Following the principle of preference transitivity Liu et al.
-(2024b), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
+(2024), we extend $P(R_{i+1}\succ R_{i})$ to $P(R_{i}\succ R_{j})$ to reduce computational cost from, thereby forming a preference matrix $P_{pre}\in\mathbb{R}^{warm\times warm}$, formally:
 $P_{pre}(R_{i}\succ R_{j})=\begin{cases}1&i=j\\ \prod_{k=j}^{i-1}P_{pre}(R_{k+1}\succ R_{k})&i>j\\ 1-{P_{pre}(R_{j}\succ R_{i})}&i<j\end{cases}$ (4)
@@
 4.1 Datasets and Tasks
-Following Yao et al.
-(2024); Suzgun and Kalai (2024); Yang et al.
-(2024a), we evaluate RoT across eight tasks (detailed in Appendix A.1): (1) Mathematical Reasoning:Game of 24 Yao et al.
+To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al.
+(2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al.
+(2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity: (1) Game of 24 Yao et al.
 (2024); Xiang et al.
-(2025), Multilingual Grade School Math Shi et al.
-(2022); (2) Domain Knowledge: Python Puzzles Schuster et al.
-(2021), Geometric Shapes, Multi-Step Arithmetic Two, Word Sorting, and Checkmate-in-One Srivastava et al.
-(2023); (3) Creativity: Shakespearean Sonnet Writing Suzgun and Kalai (2024).
+(2025) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24  [^2] https://www.4nums.com/game/ .
+We use 1000 samples drawn from a mix of both datasets for robust large-scale evaluation.
+(2) BIG-Bench Suzgun et al.
+(2023); Srivastava et al.
+(2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting and Checkmate-in-One from BIG-Bench suite  [^3] https://huggingface.co/datasets/google/bigbench Srivastava et al.
+(2023).
+(3) Python Puzzles Schuster et al.
+(2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels  [^4] https://github.com/microsoft/PythonProgrammingPuzzles.
+(4) Multilingual Grade School Math (MGSM) Shi et al.
+(2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al.
+(2021), featuring translations of a subset of examples into ten diverse languages [^5] https://github.com/google-research/url-nlp/tree/main/mgsm.
+(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three  [^6] https://huggingface.co/datasets/turingmachine/meta-prompting or five  [^7] https://github.com/iljones00/Shakespearean-Sonnets-GPT specified words verbatim.
 4.2 Baselines
@@
 4.3 Experiment Setup
-To ensure fair comparisons with previous methods, following Liu et al.
-(2024a); Yao et al.
+To ensure fair comparisons with previous methods, following Yao et al.
 (2024); Zhang et al.
 ; Suzgun and Kalai (2024), we utilize two LLMs: GPT-3.5-turbo and GPT-4, as the foundational model via OpenAI API for our RoT experiments including both main experiments and ablation study.
+For candidate generation, we use a temperature of 0.7, encouraging diversity in proposed solution.
+For instantiating reasoning steps, we adopt a lower temperature of 0.1, promoting more deterministic and stable logical progression.
 For the $warmup$ hyperparameter, we experimented with values of $\{1,3,5,10\}$, accessed in batches for reverse reasoning warm-up through OpenAI API.
@@
 (2023), offline deployed on a single NVIDIA GeForce RTX 4090 GPU.
+Knowledge Boundary and Cognitive Preference Consistency.
+As described in Section 3.3, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$.
+We first evaluate GPT-3.5-turbo & GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^9] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot.
+and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
 4.4 Evaluation Metrics
@@
 Table 1: Comparison of RoT with baselines across 1-shot and 2-shot settings for reasoning accuracy.
-Knowledge Boundary and Cognitive Preference Consistency.
-As described in Section 3.3, we trigger knowledge boundary using $S_{cog}$ by comparing $\texttt{sim}(P_{task},\mathcal{M_{LLM}}(D,P_{r})_{task})$ with a threshold $\delta=0.7$.
-We first evaluate GPT-3.5-turbo & GPT-4 on experimental tasks to distinguish between known and unknown domains in both 1-shot and 2-shot settings [^3] To achieve better cost savings and fairness, we set the few-shot for all methods to 1-shot and 2-shot.
-and subsequently conduct human evaluations of average cognitive preference consistency $\textit{Con}_{cog}$ based on stylistic and grammatical norms on $P^{*}$ across various tasks, with three professional annotators validating cognitive preferences on a 1–5 scale.
 Reasoning Accuracy.
 Following Suzgun and Kalai (2024); Yao et al.
-(2024), we introduce a LLM444we select the latest openai-o1 as judger, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong).
+(2024), we introduce a LLM101010we select the latest openai-o1 as judger during the development of our work, https://openai.com/index/introducing-openai-o1-preview/ to validate the final logical reasoning against gold results (Correct/Wrong).
 We then compute logical accuracy $Acc_{logic}$ for each logical task by tallying the number of correct responses.
@@
 5.1 Knowledge Boundary and Cognitive Preference Consistency
-Task 1-shot 2-shot
-Kno. Unkno. Kno. Unkno.
-Game of 24
-MGSM (avg)
-Multi-Step Arithmetic
-WordSorting
-Python Puzzles
-Geometric Shapes
-Checkmate-in-One
-Sonnet Writing
-Table 2: Knowledge Boundary under Reversal Reasoning for GPT-3.5-turbo and GPT-4 with identical results. Kno. for Known, Unkno. for Unknown.
 Knowledge Boundary.
@@
 (2022).
-5.2 Reasoning Accuracy and Efficiency
-(1) RoT can activate LLMs’ reasoning accuracy.
-As shown in Table 1, RoT consistently outperforms all baselines across various reasoning tasks, with particularly notable improvements observed in GPT-4.
-Specifically, compared with the best BoT, RoT achieves significant gains in tasks such as Game of 24 (+17.15% in 1-shot and +17.08% in 2-shot), Geometric Shapes (+4.87% in 1-shot and +4.96% in 2-shot), and Checkmate-in-One (+4.71% in 1-shot and +4.19% in 2-shot), demonstrating our flexibility of leveraging cognitive preference in LLM to activate logic capabilities through reverse reasoning (case study detailed in appendix B).
-For GPT-3.5-turbo, RoT also demonstrates substantial improvements, such as Game of 24 (+10.84% in 1-shot and +16.29% in 2-shot) and Geometric Shapes (+6.38% in 1-shot and +5.22% in 2-shot), further emphasizing its versatility in activating logic through cognitive preference management.
-Figure 3: Inference time comparison
+Task 1-shot 2-shot
+Kno. Unkno. Kno. Unkno.
+Game of 24
+MGSM (avg)
+Multi-Step Arithmetic
+WordSorting
+Python Puzzles
+Geometric Shapes
+Checkmate-in-One
+Sonnet Writing
+Table 2: Knowledge Boundary under Reversal Reasoning for GPT-3.5-turbo and GPT-4 with identical results. Kno. for Known, Unkno. for Unknown.
 Ablation study Game of 24 Geometric Shapes Multi-Step Arithmetic Word Sorting Checkmate-in-One Python Puzzles×$\times$ MGSM×$\times$ Sonnet Writing
@@
 Table 2: Ablation study of RoT across various tasks in 1-shot and 2-shot settings. indicates known tasks in both 1-shot and 2-shot; $\times$ indicates unknown tasks in 1-shot.
+5.2 Reasoning Accuracy and Efficiency
+(1) RoT can activate LLMs’ reasoning accuracy.
+As shown in Table 1, RoT consistently outperforms all baselines across various reasoning tasks, with particularly notable improvements observed in GPT-4.
+Specifically, compared with the best BoT, RoT achieves significant gains in tasks such as Game of 24 (+17.15% in 1-shot and +17.08% in 2-shot), Geometric Shapes (+4.87% in 1-shot and +4.96% in 2-shot), and Checkmate-in-One (+4.71% in 1-shot and +4.19% in 2-shot), demonstrating our flexibility of leveraging cognitive preference in LLM to activate logic capabilities through reverse reasoning (case study detailed in appendix B).
+For GPT-3.5-turbo, RoT also demonstrates substantial improvements, such as Game of 24 (+10.84% in 1-shot and +16.29% in 2-shot) and Geometric Shapes (+6.38% in 1-shot and +5.22% in 2-shot), further emphasizing its versatility in activating logic through cognitive preference management.
+Figure 3: Inference time comparison, measured as the average duration from inference start to evaluation end, including all steps.
 (2) RoT demonstrates better tradeoff between reasoning accuracy and efficiency across baselines.
@@
 we partially and effectively mitigates this issue through the integration of Cognitive Preference Manager (CPM) and two-shot learning.
-In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks.
+In future work, we aim to extend RoT’s capabilities by incorporating In-Context Learning (ICL) or, alternatively, adopting Auto-prompt based demonstration selection strategies Zhang et al.
+(2023); Jin and Lu (2024), which will allow for greater flexibility in adapting to varied contexts and improve its performance on more complex reasoning tasks.
 Furthermore, we believe that utilizing Reversal of Thought in teacher-student model distillation, which could further amplify the practical value of our approach.
+Finally, we acknowledge the challenges posed by the inherent instability of large language models.
+Future research may benefit from investigating robust candidate solutions—such as hybrid scoring mechanisms Wang et al.
+(2024b); Minh et al.
+(2024); Wang et al.
+(2024c)—that can be integrated with RoT to further enhance its reliability and overall performance through joint optimization.
 Ethical Considerations
@@
 Appendix A Experimental Details
-A.1 Datasets and Tasks
-To comprehensively validate our Reverse of Thought (RoT), following ToT Yao et al.
-(2024), meta-prompting Suzgun and Kalai (2024) and BoT Yang et al.
-(2024a), we assess it with baselines across a broad range of eight tasks across five logical benchmarks that encompass mathematical and algorithmic reasoning, domain-specific knowledge, and literary creativity:(1) Game of 24 Yao et al.
-(2024); Xiang et al.
-(2025) challenges LLMs to create a mathematical expression utilizing each of four given numbers exactly once to achieve 24666https://www.4nums.com/game/ .
-(2) BIG-Bench Suzgun et al.
-(2023); Srivastava et al.
-(2023) involves Geometric Shapes, Multi-Step Arithmetic Two and Word Sorting and Checkmate-in-One from BIG-Bench suite  [^7] https://huggingface.co/datasets/google/bigbench Srivastava et al.
-(2023); (3) Python Puzzles Schuster et al.
-(2021) comprises a collection of challenging programming puzzles crafted in Python, covering various difficulty levels [^8] https://github.com/microsoft/PythonProgrammingPuzzles;(4) Multilingual Grade School Math (MGSM) Shi et al.
-(2022) is a multilingual adaptation of the GSM8K dataset Cobbe et al.
-(2021), featuring translations of a subset of examples into ten diverse languages [^9] https://github.com/google-research/url-nlp/tree/main/mgsm .(5) Shakespearean Sonnet Writing Suzgun and Kalai (2024) require LLMs to compose with the rhyme scheme "ABAB CDCD EFEF GG" while incorporating three  [^10] https://huggingface.co/datasets/turingmachine/meta-prompting or five  [^11] https://github.com/iljones00/Shakespearean-Sonnets-GPT specified words verbatim.
-A.2 Prompt for Reverse Reasoning
+A.1 Prompt for Reverse Reasoning
 Prompt for Reverse Reasoning Instruction
@@
 Figure 4: Prompt for Reverse Reasoning
-A.3 Prompt for CPM (Known/Unknown)
+A.2 Prompt for CPM (Known/Unknown)
 Prompt for CPM (Known/Unknown) Instruction
@@
 If no solution exists, output the string "No feasible solution exists."
+Figure 6: $P_{final}$ Case study for Game of 24
 B.1.2 Instantiation Case Study for Game of 24 Using $P_{final}$
@@
   Output: 4 * (9 + (10 - 13))=24
+Figure 7: Instantiation Case Study for Game of 24 Using $P_{final}$
 B.2 Case Study for Unknown Task (MGSM)
@@
 Output: A single integer or float number that is the result of the operations described in the scenario.
+Figure 8: $P_{final}$ Case study for MGSM
 B.2.2 Instantiation Case Study for MGSM Using $P_{final}$
@@
   Output: 145 minutes
-Appendix C Instructions for Human Evaluation.
-To evaluate the cognitive preference consistency $Con_{cog}$ between LLMs-taste prompt for task cognitions across different tasks, annotators adhere to the following guidelines:
+Figure 9: Instantiation Case Study for MGSM Using $P_{final}$
+Appendix C Instructions for Human Evaluation
+To evaluate the cognitive preference consistency $Con_{cog}$ between LLMs-taste prompt for task cognitions across different tasks, three professional annotators adhere to the following guidelines:
 Criteria.
@@
 A score of 0-1 reflects complete inconsistency.
+Prior to the formal evaluation, several pre-scored sample responses covering every score band will be provided.
+These examples serve to calibrate annotators’ judgments and to align their understanding of the rubric.